If sinAsinB -cosAcosB = 1 show that tanA+tanB=0

Grade:11

3 Answers

ng29
8 years ago
multiply given given equation by  -1
we get cos(A+B)= (-1)   it means that A+B = pi
now tanA+tanB =0  we have to prove it 
sinA/cosA+sinB/cosB= 0  TAKING LCM  . we get sin(A+B)=0 it means we had to prove that sin(A+B)=0
and we have alredy concluded that A+B=pi  so   sin(pi) = 0
 
approve if useful
grenade
8 years ago
use the identity coa(a+b)
then you will get that 
a+b = pie
a= pie-b
substitute this value in the tan term and you will get the answer
grenade
8 years ago
approve if useful